An object with mass m is initially at rest at the origin z-0. At time t-0 it starts to accelerate with a changing acceleration along the +z direction. At time t = T it is at the point z=:T and its speed is measured as e(T) = vr 6. Find the for tr and τ in terms of F. m, and T
Answer #1

Work done by force = Change in kinetic energy
Initial kinetic energy of the object = 0 as the object starts from rest at t = 0

At time T the kinetic energy of the object = 1/2 m vT2

Work done by force = 1/2mvT2

Also, work done by the force = Force *distance traveled = FxT

Net force = mass *acceleration

a = F/m

1/2aT2 = xT

xT = 1/2F/mT2 = FT2/2m

v = aT = FT/m

Option c is correct.